OpenStudy (megannicole51):

Do the sequences converge or diverge? 3+e^(-2n) i know it diverges but how do i show out my work please:)

OpenStudy (anonymous):

i think it converges to 3, as \(n\to \infty\) you have \(e^{-2n}=\frac{1}{e^{2n}}\to 0\)

OpenStudy (megannicole51):

i put it in wolfram and it says it does not converge so idkk lol

OpenStudy (anonymous):

sequence right? not a sum

OpenStudy (agent0smith):

The *terms* do converge to 3, the sum doesn't, since you're adding 3 endlessly.

OpenStudy (anonymous):

The sequence can be rewritten as 3 + e/(2^n). As n gets very large (approaches infinity), e/2^n is e/a very large number, so e/large number approaches 0, so the entire sequence approaches 3 + 0 = 3. It converges.

OpenStudy (anonymous):

show the wolfram link, maybe i am wrong but i don't think so the sum does not converge, but this says "sequence"

OpenStudy (megannicole51):

oooh lol smart man @satellite73 ....i put it in a series not a sequence!

OpenStudy (anonymous):

guessed it right!

OpenStudy (megannicole51):

okay awesome thank you:)